### v1

@@ Abstract
Information Security; Large Language Model; Long-Text Generation; Factuality Evaluation; Multi-Agent System
@@ 1 Introduction
Research on long-form factuality evaluation seeks to develop techniques that can accurately evaluate the factual reliability of extended textual content, thereby enhancing the factual accuracy of LLM-generated outputs and mitigating hallucinations. Current evaluation approaches generally adopt atomic claim decomposition with knowledge-base verification. While methods such as FActScore [31] and SAFE [32] have pushed this direction forward, several critical challenges remain:
@@
  • Biases in Single-Model Evaluation Frameworks. Existing evaluation frameworks typically rely on a single model for factual verification, assuming it can consistently identify inaccuracies and weigh evidence [31, 32, 34, 35]. However, even leading LLMs exhibit substantial hallucination problems, which can produce incorrect or inconsistent judgments. As a result, single-model evaluation architectures are prone to systematic biases, potentially misrepresenting the factuality of generated content.
@@
  • We benchmark nine mainstream LLMs from seven model families on LongFact [32] and LongHalluQA, showing that larger models generally perform better, while Chinese-specific models excel on Chinese tasks. These results provide practical guidance for model selection and optimization in long-form factuality evaluation.
@@ 2.1 Factuality Evaluation of LLMs
2.1.1 Factuality Evaluation Benchmark
@@
In practice, LLMs are primarily applied to long-form generation tasks, whereas existing factuality benchmarks predominantly focus on short-form, manually constructed QA tasks. Benchmarks such as TruthfulQA [41], HaluEval [42], PopQA [43], FreshQA [44], and LLM-Oasis [45] assess factual consistency in English short-form QA settings. FactScore [31] was an early attempt to evaluate factuality in long-form generation but focused only on simple biographical QA. LongFact [32] expanded coverage to 38 human-curated topics across diverse domains, while FactBench [33] extracted hallucination-inducing prompts from real conversations. Nevertheless, these benchmarks remain limited to English.
In the Chinese context, HalluQA [36] targets hallucination evaluation with 450 short prompts across misleading and knowledge-based categories. ChineseSimpleQA [37] systematically evaluates models on short factual questions, including culturally specific content. However, both datasets are restricted to short-form QA, underscoring the urgent need for high-quality Chinese benchmarks for evaluating factuality in long-form generation tasks.
2.1.2 Factuality Evaluation Method
@@
Existing factuality evaluation methods for LLMs primarily follow two technical directions. One approach builds automated verifiers using advanced models such as GPT-4o [46], leveraging their strong reasoning and anomaly detection capabilities. The other approach employs retrieval-augmented generation (RAG) [47], which retrieves up-to-date knowledge to compensate for the static nature of model parameters.
Recently, long-form factuality evaluation has gained increasing attention due to the structural complexity and partial correctness of extended outputs. FactScore [31] and SAFE [32] decompose text into atomic claims for fine-grained RAG-based verification. VeriScore [34] filters unverifiable claims according to predefined criteria, and FIRE [35] introduces confidence-based retrieval to reduce computational costs.
However, most existing methods depend on a single verifier, which makes them vulnerable to systematic errors when the model hallucinates or encounters knowledge gaps, posing an open challenge that remains unresolved.
2.1.3 Factuality Evaluation Metrics
@@
Traditional evaluation paradigms rely on binary judgments at the atomic claim level and aggregate metrics such as accuracy and $F_{1}$ at the text level. FactScore [31] introduces a long-form evaluation framework by computing the proportion of factually correct atomic claims, but it does not incorporate a recall component. SAFE [32] addresses this limitation by introducing a hyperparameter K to define recall; however, its fixed value restricts adaptability across different domains and content types. FactBench [33] further accounts for claim verifiability and penalizes irrelevant content, offering a more nuanced evaluation of long-form outputs. However, all of these metrics treat facts equally, failing to account for their varying importance and relevance, which reduces sensitivity to critical errors in high-stakes applications.
@@ 2.2 Factual Verification
Factual verification has evolved from traditional automated pipelines to LLM-powered systems. Classic systems follow a three-step process: extracting verifiable claims, retrieving evidence from sources (like Wikipedia), and verifying claims using binary or multi-class classification models [48, 31, 34].
With the rise of LLMs and retrieval-augmented generation (RAG), recent systems offer stronger end-to-end performance [49]. Models like RARR [50] prompt LLMs to generate queries, retrieve evidence, and make factual judgments. Other works integrate tool use [32, 51], enabling access to external databases for dynamic knowledge updates. Multimodal extensions such as RAGAR [52] incorporate both textual and visual evidence through structured retrieval strategies like CoRAG and ToRAG.
More recently, agent-based approaches have emerged. ReAct-based agents [53] dynamically coordinate retrieval and reasoning, while multi-agent debate frameworks [54] use iterative evidence exchange to improve decision quality. These methods signal a shift from static rule-based pipelines to more adaptive, collaborative verification systems.
@@ 4.2.2 Jury: Fact Verification via Retrieval and Multi-Agent Debate
Rule 1, Autonomous Retrieval and Free Debate: In the first round, each agent may autonomously decide, based on its confidence, whether to invoke external retrieval before speaking, so as to balance knowledge utilization and retrieval cost. In the second round, agents directly deliver their statements while also engaging in discussion with others by drawing on peers’ suggestions or correcting mistakes.
Rule 2, Mandatory Retrieval and Evidence-Based Debate: In the first round, retrieval is mandatory before speaking, ensuring that agents compensate for possible knowledge gaps and mitigate overconfidence, a phenomenon where they assign overly high confidence to themselves. In the second round, agents debate with others based on multiple retrieved references, making the discussion more evidence-grounded and reliable.
Rule 3, Dynamic Retrieval and Adaptive Debate: In the first round, each agent may autonomously retrieve information before speaking. The debate process in later rounds is then adjusted according to the jury’s consensus: if consensus is reached in the first round, the results are directly output to avoid redundant costs; if not, retrieval becomes mandatory before the second round, so that authoritative references can help resolve conflicts and guide the jury toward agreement.
Agent Role-Playing
@@ 4.2.3 Judge: Determining the Final Result
For a long-text response $a_{i}$ generated by the evaluated model, along with its corresponding T atomic claims $\{c_{i,j}\}^{T}_{j=1}$, the Judge agent integrates the final evaluation results of each atomic claim $\{p_{i,j}\}^{T}_{j=1}$, to calculate the score $s_{i}$ of the response. This process can be expressed as:
@@
The overall score $s$ of the evaluated model on the entire dataset is defined as the arithmetic mean of the scores $s_{i}$ of all long-text responses, which can be formulated as $s=\frac{1}{|D|}\sum_{i=1}^{|D|}s_{i}$, where $|D|$ denotes the total number of samples in the dataset. The specific calculation method will be elaborated in the next chapter.
@@ 5.1 Fact Importance Hierarchy Model
$\{g_{i,k}\}^{K_{gold}}_{k=1}=\mathrm{set}(\{[c_{i,j,k}]^{K_{j}}_{k=1}\}^{G}_{j=1}),$ (16)
@@ 5.2 Weighted Evaluation Metrics
Given a question $q_{i}$, we construct the pyramid model $P_{i,G}$ following the aforementioned procedure and obtain the golden set of atomic claims $\{g_{i,k}\}^{K_{gold}}_{k=1}$. For a long-text response $a_{i}$ to be evaluated, we calculate the weight $\omega_{i,j}$ of each decomposed atomic claim $c_{i,j}$ according to the pyramid model and define the following weighted evaluation metrics.
@@ 5.2.1 Weighted Precision
$Prec_{w}(a_{i})=$ $\frac{\sum_{j=1}^{|S|}\omega_{i,j}\,c^{T}_{i,j}}{\sum_{j=1}^{|S|}\omega_{i,j}\,c^{T}_{i,j}+\sum_{j=1}^{|N|}\omega_{i,j}\,c^{F}_{i,j}}$ (17)
$=$ $\frac{\sum_{j=1}^{|S|}\omega_{i,j}\,c^{T}_{i,j}}{\sum_{j=1}^{|S|+|N|}\omega_{i,j}\,c_{i,j}}.$
@@ 5.2.3 Weighted F1-Score
$F_{1}@\gamma(a_{i})=\begin{cases}\frac{2\cdot Prec_{w}(a_{i})\cdot R_{w}@\gamma(a_{i})}{Prec_{w}(a_{i})+R_{w}@\gamma(a_{i})},&\text{if }|S|>0,\\ 0,&\text{if }|S|=0.\end{cases}$ (19)
@@ 6.1.1 Datasets
All five datasets contain multiple long-text factual QA pairs, along with the corresponding atomic claims decomposed from the long responses and their binary factuality labels. Detailed statistics are shown in Table 2.
@@ 6.1.2 Baselines
(1) SAFE [32]: Utilizes LLM-based agents for long-text factuality evaluation, outperforming crowdsourced human annotators by reaching 72% agreement and prevailing in 76% of 100 randomly sampled disagreement cases.
@@ 6.1.3 Evaluation Metrics
$\text{Prec}=\frac{TP}{TP+FP},~~~~\text{Recall}=\frac{TP}{TP+FN},~~~~F_{1}=\frac{2\times\text{Prec}\times\text{Recall}}{\text{Prec}+\text{Recall}}.$ (20)
@@ 6.1.4 Comparative Experiments
Precision Recall F1 Precision Recall F1
Factcheck-Bench SAFE 0.89 0.83 0.86 0.70 0.79 0.74
@@
Precision Recall F1 Precision Recall F1
Factcheck-Bench SAFE 0.94 0.74 0.83 0.62 0.90 0.74
@@
Precision Recall F1 Precision Recall F1
Factcheck-Bench MAD-Fact 0.94 0.77 0.84 0.55 0.84 0.67
@@ 6.1.5 Ablation Study
(1)w/o Role-Playing: Removing the diversified role configurations (i.e., modifying the agent profile $conf^{m}$ in Equation (7)), such that all Evaluator agents have exactly the same role settings;
@@
(3)w/o Search: Removing the retrieval module (i.e., Equation (8)), prohibiting Evaluator agents from invoking external retrieval tools to access reference materials before their statements.
@@
Precision Recall F1 Precision Recall F1
@@
+w/o Role-Playing 0.89 0.87 0.88 0.70 0.74 0.72
+w/o Debate 0.87 0.78 0.82 0.58 0.71 0.64
+w/o Search 0.85 0.83 0.84 0.63 0.52 0.57
@@
+w/o Role-Playing 0.91 0.83 0.87 0.58 0.75 0.66
+w/o Debate 0.89 0.85 0.87 0.57 0.66 0.61
+w/o Search 0.84 0.84 0.84 0.48 0.48 0.48
@@ 6.2.1 Datasets
This study employs two long-text factuality evaluation datasets: LongFact [32] and LongHalluQA. LongFact is the first systematically constructed cross-domain benchmark for long-text factuality evaluation, encompassing 38 manually curated topics and 2,280 factual questions, where models are required to generate long-form textual answers. Following the setup of SAFE [32], we apply stratified sampling to select 250 instances from LongFact with a balanced topic distribution as our test set. LongHalluQA, newly introduced in this paper, is a Chinese benchmark for long-text factuality evaluation, with details provided in Section 3. From this dataset, we sampled 250 high-quality instances for evaluation.
@@ 6.2.2 Evaluated Models
  • Closed-source models: OpenAI series: GPT-3.5-Turbo, GPT-4-Turbo; Claude series: Claude-3-Sonnet; Doubao series: Doubao-1.5-Pro-32k; Kimi series: Moonshot-V1-8k.
@@ 6.2.4 Results Analysis
As shown in Figure 11 and Table 7, models with larger parameter sizes and more recent releases generally demonstrate better factual consistency in long-text scenarios. For instance, GPT-4-Turbo outperforms GPT-3.5-Turbo. Under the two selected $\gamma$ values, the best-performing model in terms of factuality is GPT-4-Turbo, consistent with the findings reported by SAFE [32]. The performance gap between the domestic model Doubao-1.5-Pro and GPT-4-Turbo is minimal, only about 1–2 percentage points, highlighting the strong potential and advantages of domestic models. GPT-3.5-Turbo, as a relatively older model, is surpassed by most newer models. It is noteworthy that Claude-3.7-Sonnet and QwQ-32B, both widely regarded as strong and recently released models, did not achieve high rankings. This is primarily because Claude-3.7-Sonnet has a relatively low factual recall score (0.501 under $\gamma=0.8$, ranking 8th among the selected models), while QwQ-32B has a relatively low factual precision score (0.763, ranking 9th). These results indicate that strong reasoning or coding capabilities do not necessarily guarantee superior factuality.
@@
As shown in Figure 12 and Table 8, domestic models exhibit notable advantages in factuality evaluation for Chinese long-text responses. Under the two selected $\gamma$ values, the top three models in terms of factuality are QwQ-32B, Doubao-1.5-Pro, and DeepSeek-V3, all domestic models. Among them, QwQ-32B and Doubao-1.5-Pro clearly outperform the others, exceeding the third-ranked DeepSeek-V3 by nearly 10 percentage points. In contrast, GPT-4-Turbo performs worse on the Chinese dataset than on the English dataset, ranking 5th among the selected models when $\gamma=1.0$ and 6th when $\gamma=0.8$, indicating a notable issue of cultural bias [66] and an imbalance in multilingual capabilities.
@@
To verify the effectiveness of the weighted evaluation metric based on the hierarchical importance of facts, we established a review panel consisting of two individuals with undergraduate degrees. From the LongFact dataset, 50 questions were sampled, and the outputs of different models were manually rated. The correlation between the human ratings and the weighted F1 score ($\gamma=0.8$) was then calculated, as shown in Figure 13. The results indicate that the Pearson correlation coefficient reached a statistically significant value of $r=0.701$ ($p=0.036$), demonstrating that the proposed weighted evaluation metric aligns well with human judgments in assessing the factuality of long-text responses.

### v2

@@ Abstract
Information Security; Large Language Model; Long-Form Text Generation; Factuality Evaluation; Multi-Agent System
[…]
@@ 1 Introduction
Research on long-form factuality evaluation seeks to develop techniques that can accurately evaluate the factual reliability of extended textual content, thereby enhancing the factual accuracy of outputs generated by LLMs and mitigating hallucinations. Current approaches generally adopt atomic claim decomposition with knowledge-base verification. While methods such as FActScore [31] and SAFE [32] have pushed this direction forward, several critical challenges remain:
@@
  • Biases in Single-Model Evaluation Frameworks. Existing evaluation frameworks typically rely on a single model for factual verification, assuming it can consistently identify inaccuracies and weigh evidence [31, 32, 34, 35]. However, even leading LLMs exhibit substantial hallucination problems [Hughes_Vectara_Hallucination_Leaderboard_2023], which can produce incorrect or inconsistent judgments. As a result, single-model evaluation architectures are prone to systematic biases, potentially misrepresenting the factuality of generated content.
@@
  • We benchmark 9 mainstream LLMs from 7 model families on LongFact [32] and LongHalluQA, showing that larger models generally perform better, while Chinese-specific models excel on Chinese tasks. These results provide practical guidance for model selection and optimization in long-form factuality evaluation.
@@ 2.1 Factuality Evaluation of LLMs
Evaluating the factuality of LLMs has been explored through diverse benchmarks, methods, and metrics. This section reviews recent progress along these three dimensions, highlighting current limitations in long-form generation scenarios.
@@
Factuality Evaluation Benchmark. In practice, LLMs are primarily applied to long-form generation tasks, whereas existing factuality benchmarks predominantly focus on short-form, manually constructed QA tasks. Benchmarks such as TruthfulQA [41], HaluEval [42], PopQA [43], FreshQA [44], and LLM-Oasis [45] assess factual consistency in English short-form QA settings. FactScore [31] was an early attempt to evaluate factuality in long-form generation but focused only on simple biographical QA. LongFact [32] expanded coverage to 38 human-curated topics across diverse domains, while FactBench [33] extracted hallucination-inducing prompts from real conversations. Nevertheless, these benchmarks remain limited to English. In the Chinese context, HalluQA [36] targets hallucination evaluation with 450 short prompts across misleading and knowledge-based categories. ChineseSimpleQA [37] systematically evaluates models on short factual questions, including culturally specific content. However, both datasets are restricted to short-form QA, underscoring the urgent need for high-quality Chinese benchmarks for evaluating factuality in long-form generation tasks.
@@
Factuality Evaluation Method. Existing factuality evaluation methods for LLMs primarily follow two technical directions. One approach builds automated verifiers using advanced models such as GPT-4o [46], leveraging their strong reasoning and anomaly detection capabilities. The other approach employs retrieval-augmented generation (RAG) [47], which retrieves up-to-date knowledge to compensate for the static nature of model parameters. Recently, long-form factuality evaluation has gained increasing attention due to the structural complexity and partial correctness of extended outputs. FactScore [31] and SAFE [32] decompose text into atomic claims for fine-grained RAG-based verification. VeriScore [34] filters unverifiable claims according to predefined criteria, and FIRE [35] introduces confidence-based retrieval to reduce computational costs. However, most existing methods depend on a single verifier, which makes them vulnerable to systematic errors when the model hallucinates or encounters knowledge gaps, posing an open challenge that remains unresolved.
@@
Factuality Evaluation Metrics. Traditional evaluation paradigms rely on binary judgments at the atomic claim level and aggregate metrics such as accuracy and $F_{1}$ at the text level. FactScore [31] introduces a long-form evaluation framework by computing the proportion of factually correct atomic claims, but it does not incorporate a recall component. SAFE [32] addresses this limitation by introducing a hyperparameter K to define recall; however, its fixed value restricts adaptability across different domains and content types. FactBench [33] further accounts for claim verifiability and penalizes irrelevant content, offering a more nuanced evaluation of long-form outputs. However, all of these metrics treat facts equally, failing to account for their varying importance and relevance, which reduces sensitivity to critical errors in high-stakes applications.
@@ 2.2 Factual Verification
Factual verification has evolved from traditional automated pipelines to LLM-powered systems. Classic systems follow a three-step process: extracting verifiable claims, retrieving evidence from sources (like Wikipedia), and verifying claims using binary or multi-class classification models [48, 31, 34]. With the rise of LLMs and retrieval-augmented generation (RAG), recent systems offer stronger end-to-end performance [49]. Models like RARR [50] prompt LLMs to generate queries, retrieve evidence, and make factual judgments. Other works integrate tool use [32, 51], enabling access to external databases for dynamic knowledge updates. Multimodal extensions such as RAGAR [52] incorporate both textual and visual evidence through structured retrieval strategies like CoRAG and ToRAG. More recently, agent-based approaches have emerged. ReAct-based agents [53] dynamically coordinate retrieval and reasoning, while multi-agent debate frameworks [54] use iterative evidence exchange to improve decision quality. These methods signal a shift from static rule-based pipelines to more adaptive, collaborative verification systems.
@@ 4.2.2 Jury: Fact Verification via Retrieval and Multi-Agent Debate
Rule 1: Autonomous Retrieval and Free Debate. In the first round, each agent may autonomously decide, based on its confidence, whether to invoke external retrieval before speaking, so as to balance knowledge utilization and retrieval cost. In the second round, agents directly deliver their statements while also engaging in discussion with others by drawing on peers’ suggestions or correcting mistakes.
Rule 2: Mandatory Retrieval and Evidence-Based Debate. In the first round, retrieval is mandatory before speaking, ensuring that agents compensate for possible knowledge gaps and mitigate overconfidence, a phenomenon where they assign overly high confidence to themselves. In the second round, agents debate with others based on multiple retrieved references, making the discussion more evidence-grounded and reliable.
Rule 3: Dynamic Retrieval and Adaptive Debate. In the first round, each agent may autonomously retrieve information before speaking. The debate process in later rounds is then adjusted according to the jury’s consensus: if consensus is reached in the first round, the results are directly output to avoid redundant costs; if not, retrieval becomes mandatory before the second round, so that authoritative references can help resolve conflicts and guide the jury toward agreement.
@@ 4.2.3 Judge: Determining the Final Result
For a long-form response $a_{i}$ generated by the evaluated model, along with its corresponding T atomic claims $\{c_{i,j}\}^{T}_{j=1}$, the Judge agent integrates the final evaluation results of each atomic claim $\{p_{i,j}\}^{T}_{j=1}$, to calculate the score $s_{i}$ of the response. This process can be expressed as:
@@
The overall score $s$ of the evaluated model on the entire dataset is defined as the arithmetic mean of the scores $s_{i}$ of all long-form responses, which can be formulated as $s=\frac{1}{|D|}\sum_{i=1}^{|D|}s_{i}$, where $|D|$ denotes the total number of samples in the dataset. The specific calculation method will be elaborated in the next chapter.
@@ 5.1 Fact Importance Hierarchy Model
$\{g_{i,k}\}^{K_{gold}}_{k=1}=\mathrm{set}\left(\{[c_{i,j,k}]^{K_{j}}_{k=1}\}^{G}_{j=1}\right),$ (16)
@@ 5.2 Weighted Evaluation Metrics
Given a question $q_{i}$, we construct the pyramid model $P_{i,G}$ following the aforementioned procedure and obtain the golden set of atomic claims $\{g_{i,k}\}^{K_{gold}}_{k=1}$. For a long-form response $a_{i}$ to be evaluated, we calculate the weight $\omega_{i,j}$ of each decomposed atomic claim $c_{i,j}$ according to the pyramid model and define the following weighted evaluation metrics.
@@ 5.2.1 Weighted Precision
$aligned Prec_{w}(a_{i})=&\frac{\sum_{j=1}^{|S|}\omega_{i,j}\,c^{T}_{i,j}}{\sum_{j=1}^{|S|}\omega_{i,j}\,c^{T}_{i,j}+\sum_{j=1}^{|N|}\omega_{i,j}\,c^{F}_{i,j}}\\ =&\frac{\sum_{j=1}^{|S|}\omega_{i,j}\,c^{T}_{i,j}}{\sum_{j=1}^{|S|+|N|}\omega_{i,j}\,c_{i,j}}.$ (17)
@@ 5.2.3 Weighted F1-Score
[…]
@@ 6.1.1 Datasets
All five datasets contain multiple long-form factual QA pairs, along with the corresponding atomic claims decomposed from the long responses and their binary factuality labels. Detailed statistics are shown in Table 2.
@@ 6.1.2 Baselines
(1) SAFE [32]: Utilizes LLM-based agents for long-form factuality evaluation, outperforming crowdsourced human annotators by reaching 72% agreement and prevailing in 76% of 100 randomly sampled disagreement cases.
@@ 6.1.3 Evaluation Metrics
[…]
@@ 6.1.4 Comparative Experiments
Prec Recall F1 Prec Recall F1
FactcheckBench SAFE 0.89 0.83 0.86 0.70 0.79 0.74
@@
Prec Recall F1 Prec Recall F1
FactcheckBench SAFE 0.94 0.74 0.83 0.62 0.90 0.74
@@
Prec Recall F1 Prec Recall F1
FactcheckBench MAD-Fact 0.94 0.77 0.84 0.55 0.84 0.67
@@ 6.1.5 Ablation Study
[…]
@@
[…]
@@
Prec Recall F1 Prec Recall F1
@@
w/o Role-Playing 0.89 0.87 0.88 0.70 0.74 0.72
w/o Debate 0.87 0.78 0.82 0.58 0.71 0.64
w/o Search 0.85 0.83 0.84 0.63 0.52 0.57
@@
w/o Role-Playing 0.91 0.83 0.87 0.58 0.75 0.66
w/o Debate 0.89 0.85 0.87 0.57 0.66 0.61
w/o Search 0.84 0.84 0.84 0.48 0.48 0.48
@@ 6.2.1 Datasets
This study employs two long-form factuality evaluation datasets: LongFact [32] and LongHalluQA. LongFact is the first systematically constructed cross-domain benchmark for long-form factuality evaluation, encompassing 38 manually curated topics and 2,280 factual questions, where models are required to generate long-form textual answers. Following the setup of SAFE [32], we apply stratified sampling to select 250 instances from LongFact with a balanced topic distribution as our test set. LongHalluQA, newly introduced in this paper, is a Chinese benchmark for long-form factuality evaluation, with details provided in Section 3. From this dataset, we sampled 250 high-quality instances for evaluation.
@@ 6.2.2 Evaluated Models
  • Closed-source models: Claude series: Claude-3-Sonnet; OpenAI series: GPT-3.5-Turbo, GPT-4-Turbo; Doubao series: Doubao-1.5-Pro-32k; Kimi series: Moonshot-V1-8k.
@@ 6.2.4 Results Analysis
As shown in Figure 11 and Table 7, models with larger parameter sizes and more recent releases generally demonstrate better factual consistency in long-form scenarios. For instance, GPT-4-Turbo outperforms GPT-3.5-Turbo. Under the two selected $\gamma$ values, the best-performing model in terms of factuality is GPT-4-Turbo, consistent with the findings reported by SAFE [32]. The performance gap between the domestic model Doubao-1.5-Pro and GPT-4-Turbo is minimal, only about 1–2 percentage points, highlighting the strong potential and advantages of domestic models. GPT-3.5-Turbo, as a relatively older model, is surpassed by most newer models. It is noteworthy that Claude-3.7-Sonnet and QwQ-32B, both widely regarded as strong and recently released models, did not achieve high rankings. This is primarily because Claude-3.7-Sonnet has a relatively low factual recall score (0.501 under $\gamma=0.8$, ranking 8th among the selected models), while QwQ-32B has a relatively low factual precision score (0.763, ranking 9th). These results indicate that strong reasoning or coding capabilities do not necessarily guarantee superior factuality.
@@
As shown in Figure 12 and Table 8, domestic models exhibit notable advantages in factuality evaluation for Chinese long-form responses. Under the two selected $\gamma$ values, the top three models in terms of factuality are QwQ-32B, Doubao-1.5-Pro, and DeepSeek-V3, all domestic models. Among them, QwQ-32B and Doubao-1.5-Pro clearly outperform the others, exceeding the third-ranked DeepSeek-V3 by nearly 10 percentage points. In contrast, GPT-4-Turbo performs worse on the Chinese dataset than on the English dataset, ranking 5th among the selected models when $\gamma=1.0$ and 6th when $\gamma=0.8$, indicating a notable issue of cultural bias [66] and an imbalance in multilingual capabilities.
@@
To verify the effectiveness of the weighted evaluation metric based on the hierarchical importance of facts, we established a review panel consisting of two individuals with undergraduate degrees. From the LongFact dataset, 50 questions were sampled, and the outputs of different models were manually rated. The correlation between the human ratings and the weighted $F_{1}$ score ($\gamma=0.8$) was then calculated, as shown in Figure 13. The results indicate that the Pearson correlation coefficient reached a statistically significant value of $r=0.701$ ($p=0.036$), demonstrating that the proposed weighted evaluation metric aligns well with human judgments in assessing the factuality of long-form responses.